# XLSR-Mamba: A Dual-Column Bidirectional State Space Model for Spoofing Attack Detection

### v1

@@ Abstract
Transformers and their variants have achieved great success in speech processing. However, their multi-head self-attention mechanism is computationally expensive. Therefore, one novel selective state space model, Mamba, has been proposed as an alternative. Building on its success in automatic speech recognition, we apply Mamba for spoofing attack detection. Mamba is well-suited for this task as it can capture the artifacts in spoofed speech signals by handling long-length sequences. However, Mamba’s performance may suffer when it is trained with limited labeled data. To mitigate this, we propose combining a new structure of Mamba based on a dual-column architecture with self-supervised learning, using the pre-trained wav2vec 2.0 model. The experiments show that our proposed approach achieves competitive results and faster inference on the ASVspoof 2021 LA and DF datasets, and on the more challenging In-the-Wild dataset, it emerges as the strongest candidate for spoofing attack detection. The code will be publicly released in due course.
@@ I Introduction
Previous works in anti-spoofing advocated long-term features to capture the relevant artifacts for spoofing attack detection [7, 8, 9]. Adding value to these works, recent research shows that the cues discriminating spoofed speech from bonafide ones can be well apprehended by local and global features [10, 11]. Local features may reveal unnatural intonation or stress, while global features often exhibit irregular rhythms or flat emotional tones. Combining information from both levels enhances detection performance.
@@
Self-supervised learning (SSL) [13] has gained attention for creating pre-trained models that generalize well with limited labeled data. Several SSL speech models, including auto-regressive predictive coding, wav2vec [14], and HuBERT [15], have demonstrated promising results in various speech-processing tasks. A recent anti-spoofing model combined the SSL model wav2vec 2.0 XLS-R (0.3B) [16, 17] with the transformer-based conformer architecture [18, 19], achieving state-of-the-art (SOTA) results in ASVspoof 2021 [20]. This success is linked to the multi-head self-attention (MHSA) [21] mechanism’s strong sequence modeling capabilities. However, empirical evidence shows that MHSA is less effective at selectively capturing relevant information for modeling long-range relationships. Therefore, we are interested in proposing a more effective method for capturing long-range feature information in speech signals. Furthermore, the self-attention mechanism faces rising computational complexity with large context windows. To address these issues, state space models (SSMs) [22] like Mamba [23] have been introduced very recently, showing promising results. The RawBMamba [24] model introduces a bidirectional Mamba in anti-spoofing. Nevertheless, the challenge lies in effectively applying Mamba for anti-spoofing tasks with pre-trained models as an alternative to transformer-based frameworks.
@@
Figure 1: Overview of the XLSR-Mamba architecture for anti-spoofing, with three different BiMamba configurations: (a) External Bidirectional Mamba (ExtBiMamba); (b) Inner Bidirectional Mamba (InnBiMamba); and (c) Dual-Column Bidirectional Mamba (DuaBiMamba); (d) XLSR-Mamba model pipeline, including XLS-R feature extraction, linear projection, BiMamba blocks, and prediction head.
@@ II Preliminaries
Structured SSMs efficiently handle long-dependent sequences with low computational and memory demands, making them suitable replacements for transformers or recurrent neural networks (RNNs) [22]. Mamba improves upon SSMs by introducing an input-dependent selection mechanism for better information filtering and a hardware-aware algorithm that scales linearly with sequence length, allowing for faster computation [23]. Mamba’s architecture combines SSM blocks with linear layers. It is simple yet achieves the SOTA performance in various long-sequence tasks [25], offering substantial computational efficiency during training and inference.
At its core, Mamba employs a linear selective SSM. In the equation below, the variables $h_{t}$, $x_{t}$, and $y_{t}$ represent the hidden state, input, and output at time $t$, respectively. The matrices $A$, $B$, and $C$ are learnable parameters corresponding to the state transition, input, and output processes.
$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$ (1)
@@
$\quad y=x\ast\mathcal{K},{\text{~{}where~{}}}\mathcal{K}=(CB,AB,\ldots,CA^{\mathcal{L}-1}B)$ (2)
However, since the matrices $A_{t}$, $B_{t}$, and $C_{t}$ depend on the input $x_{t}$ (making them selective), this convolution cannot be computed directly. Instead, it is solved using a parallel scan algorithm. A unidirectional Mamba is an SSM placed between gated linear layers. While these modifications have improved model performance by adding “selective” features, they do not change the fundamental unidirectional nature of SSMs, similar to RNNs. This unidirectionality is not a problem for training large language models, as many use an autoregressive approach. However, in non-autoregressive speech models for anti-spoofing, a module with non-causal capabilities, like attention mechanisms, is necessary. Thus, it is essential to find an effective solution to address this limitation.
@@ III-B Proposed Dual-Column Bidirectional Mamba
Both InnBiMamba and ExtBiMamba have limited adaptability for anti-spoofing tasks, as they lack mechanisms to capture the nuances of local and global features needed for spoofing attack detection. In this regard, we propose the Dual-Column Bidirectional Mamba (DuaBiMamba) to address these limitations. DuaBiMamba as shown in Fig. 1 (c) is designed with two separate Mamba columns: one column processes all forward features, while the other handles all backward features. The final outputs from each direction are fused by concatenation, providing a richer, bidirectional representation that combines local and global feature dependencies. Thereby, this dual-column based Mamba structure captures the complex patterns essential for anti-spoofing.
@@ III-C Proposed XLSR-Mamba
We adopt the XLSR-Conformer [18] structure as our baseline architecture to propose the XLSR-Mamba model as shown in Fig. 1 (d). This model leverages the pre-trained XLSR, a variant of wav2vec 2.0. With extensive SSL training, the model can extract rich speech representations for spoofing attack detection. The XLSR model consists of two primary components: a convolutional neural network front-end, which converts the 1D raw waveform into a 2D temporal-channel representation, and 24 transformer encoder layers that capture global relationships within the speech signal. The resulting speech representation has a shape of $(T\times C)$, where $T$ denotes temporal length and $C$ represents the channel dimension. After obtaining the XLSR representation, we project it to a $D$-dimensional space. These $(T\times D)$ features are then passed through the BiMamba blocks to learn higher-level representations. The BiMamba composes $L$ Mamba blocks, each containing SSM and Conv1D layers to capture local dependencies in the speech representation. Finally, the output features from the BiMamba model are used to classify the given input speech as bonafide or spoofed.
@@ IV-B Implementation Details
During training, audio segments were cropped or concatenated to around 4 seconds (64,600 samples). The Adam optimizer was used with a learning rate of $10^{-6}$ and a weight decay of $10^{-4}$ to optimize a weighted cross-entropy loss, with a batch size of 20. Final results were based on an averaged checkpoint of the top-5 validation models, with early stopping after 7 epochs without improvement. We set the embedding size of XLSR to 144 and $L$ of Mamba blocks to 12 as the default configuration.
@@ V Results and Analysis
Table I shows that DuaBiMamba performs the best on the ASVspoof 2021 LA set, achieving the lowest EER of 0.93% and min t-DCF. Its dual-column structure, with separate forward and backward Mamba layers, excels at capturing local and global dependencies, which is particularly beneficial for identifying fine-grained artifacts in the LA data. However, InnBiMamba achieves the lowest EER of 1.82% on the DF set. This may be due to its shared input and output projection layers, which create a more coherent representation suited to the consistent, compression-induced distortions in the DF dataset. ExtBiMamba, while providing separate pathways, may not capture the interaction between forward and backward dependencies as effectively as DuaBiMamba’s concatenation approach. It is also noted that the performance declines without BiMambas as shown in the first row of Table I, indicating its critical role. As DuaBiMamba is better at detecting diverse artifacts of LA and DF data collectively, we propose to use it as the backbone for the XLSR-Mamba model.
@@ V-D Evaluation on In-the-Wild dataset
We also evaluate our proposed model with the In-the-Wild dataset which is more challenging and capable of reflecting the generality of models due to similar background noise, emotional tone, and duration across spoofed and genuine instances. Table III shows that the proposed XLSR-Mamba model achieves the lowest EER of 6.71%, outperforming other SOTA single systems on this challenging dataset. XLSR-Mamba’s structure enhances its sensitivity to long-range timing and rhythm inconsistencies, which are often subtle in deepfake speech and essential for accurately distinguishing between real and spoofed samples in realistic scenarios.

### v2

@@ Abstract
Transformers and their variants have achieved great success in speech processing. However, their multi-head self-attention mechanism is computationally expensive. Therefore, one novel selective state space model, Mamba, has been proposed as an alternative. Building on its success in automatic speech recognition, we apply Mamba for spoofing attack detection. Mamba is well-suited for this task as it can capture the artifacts in spoofed speech signals by handling long-length sequences. However, Mamba’s performance may suffer when it is trained with limited labeled data. To mitigate this, we propose combining a new structure of Mamba based on a dual-column architecture with self-supervised learning, using the pre-trained wav2vec 2.0 model. The experiments show that our proposed approach achieves competitive results and faster inference on the ASVspoof 2021 LA and DF datasets, and on the more challenging In-the-Wild dataset, it emerges as the strongest candidate for spoofing attack detection. The code has been publicly released [^1] https://github.com/swagshaw/XLSR-Mamba.
@@ I Introduction
Previous works in anti-spoofing advocated long-term features to capture the relevant artifacts for spoofing attack detection [7, 8, 9, 10]. Adding value to these works, recent research shows that the cues discriminating spoofed speech from bonafide ones can be well apprehended by local and global features [11, 12]. Local features may reveal unnatural intonation or stress, while global features often exhibit irregular rhythms or flat emotional tones. Combining information from both levels enhances detection performance.
@@
Self-supervised learning (SSL) [14] has gained attention for creating pre-trained models that generalize well with limited labeled data. Several SSL speech models, including auto-regressive predictive coding, wav2vec [15], and HuBERT [16], have demonstrated promising results in various speech-processing tasks. A recent anti-spoofing model combined the SSL model wav2vec 2.0 XLS-R (0.3B) [17, 18] with the transformer-based conformer architecture [19, 20], achieving state-of-the-art (SOTA) results in ASVspoof 2021 [21]. This success is linked to the multi-head self-attention (MHSA) [22] mechanism’s strong sequence modeling capabilities. However, empirical evidence [23] shows that MHSA is less effective at selectively capturing relevant information for modeling long-range relationships. Therefore, we are interested in proposing a more effective method for capturing long-range feature information in speech signals. Furthermore, the self-attention mechanism faces rising computational complexity with large context windows. To address these issues, state space models (SSMs) [24] like Mamba [25] have been introduced very recently, showing promising results. The RawBMamba [26] introduces a bidirectional Mamba in anti-spoofing. Nevertheless, the challenge lies in effectively applying Mamba for anti-spoofing tasks with pre-trained models as an alternative to transformer-based frameworks.
@@
Figure 1: Overview of the XLSR-Mamba architecture for anti-spoofing, with three different BiMamba configurations: (a) External Bidirectional Mamba (ExtBiMamba); (b) Inner Bidirectional Mamba (InnBiMamba); and (c) Dual-Column Bidirectional Mamba (DuaBiMamba); (d) XLSR-Mamba model pipeline, including XLS-R feature extraction, linear projection, BiMamba blocks, and prediction head. The prediction head is implemented as a linear layer that maps features produced by the BiMamba blocks into whether the input speech is bonafide or spoofed.
@@ II Preliminaries
Structured SSMs [24] efficiently handle long-dependent sequences with low computational and memory demands, making them suitable replacements for transformers or recurrent neural networks (RNNs). Mamba improves upon SSMs by introducing an input-dependent selection mechanism for better information filtering and a hardware-aware algorithm that scales linearly with sequence length, allowing for faster computation [25]. Mamba’s architecture combines SSM blocks with linear layers. It is simple yet achieves the SOTA performance in various long-sequence tasks [27], offering substantial computational efficiency during training and inference.
At its core, Mamba employs a linear selective SSM. The SSM operates by mapping an input $x_{t}$ to an output $y_{t}$ through a hidden state $h_{t}$ at time $t$. The matrices $A$, $B$, and $C$ are continuous learnable parameters corresponding to the state transition, input, and output processes. We first discretize the continuous parameters to discrete counterparts $\tilde{A}$, $\tilde{B}$. This operation allows the model to process discrete-time audio signals. Then the SSM as the equation below:
$h_{t}=\tilde{A}h_{t-1}+\tilde{B}x_{t},\quad y_{t}=Ch_{t}$ (1)
@@
$\quad y=x\ast\mathcal{K},{\text{~{}where~{}}}\mathcal{K}=(C\tilde{B},C\tilde{A}\tilde{B},\ldots,C\tilde{A}^{\mathcal{L}-1}\tilde{B})$ (2)
Here, each element in $\mathcal{K}$ reflects how an input at a given time step is propagated through the state transitions to influence future outputs. This convolution kernel $\mathcal{K}$ compactly captures the temporal dynamics of the model, enabling efficient processing of long sequences—a key advantage of structured state-space models for anti-spoofing tasks. However, since the matrices depend on the input $x_{t}$ (making them selective), this convolution cannot be computed directly. Instead, it is solved using a parallel scan algorithm. A unidirectional Mamba is an SSM placed between gated linear layers. While these modifications have improved model performance by adding “selective” features, they do not change the fundamental unidirectional nature of SSMs, similar to RNNs. This unidirectionality is not a problem for training large language models, as many use an autoregressive approach. However, in non-autoregressive speech models for anti-spoofing, a module with non-causal capabilities, like attention mechanisms, is necessary. Thus, it is essential to find an effective solution to address this limitation.
@@ III-B Proposed Dual-Column Bidirectional Mamba
Both InnBiMamba and ExtBiMamba have limited adaptability for anti-spoofing tasks. In these models, the forward and backward pathways are merged either through shared input/output projections (InnBiMamba) or simple addition (ExtBiMamba), which restricts their ability to differentiate fine-grained cues—such as unnatural intonation or stress—from broader patterns like a flat emotional tone. In this regard, we propose the Dual-Column Bidirectional Mamba (DuaBiMamba) to address these limitations. DuaBiMamba as shown in Fig. 1 (c) is designed with two separate Mamba columns: one column processes all forward features, while the other handles all backward features. Their outputs are then concatenated, which fuses local details (captured via Conv1D layers within each Mamba block) and global dependencies (modeled by the selective state space mechanisms). Thereby, this dual-column based Mamba structure captures the complex patterns essential for anti-spoofing.
@@ III-C Proposed XLSR-Mamba
We adopt the XLSR-Conformer [19] structure as our baseline architecture to propose the XLSR-Mamba model as shown in Fig. 1 (d). This model leverages the pre-trained XLSR, a variant of wav2vec 2.0. With extensive SSL training, the model can extract rich speech representations for spoofing attack detection. The XLSR model consists of two primary components: a convolutional neural network front-end, which converts the 1D raw waveform into a 2D temporal-channel representation, and 24 transformer encoder layers that capture global relationships within the speech signal. The resulting speech representation is a $T\times 1024$ matrix, where $T$ denotes temporal length and 1024 is the number of channels. After obtaining the XLSR representation, we project it to a $D$-dimensional space. These $(T\times D)$ features are then passed through the BiMamba blocks to learn higher-level representations. The BiMamba composes $N$ Mamba blocks, each containing SSM and Conv1D layers to capture local dependencies in the speech representation. Finally, the output features from the BiMamba model are used to classify the given input speech as bonafide or spoofed.
@@ IV-B Implementation Details
During training, audio segments were cropped or concatenated to around 4 seconds (64,600 samples). The Adam optimizer was used with a learning rate of $10^{-6}$ and a weight decay of $10^{-4}$ to optimize a weighted cross-entropy loss, with a batch size of 20. Final results were based on an averaged checkpoint of the top-5 validation models, with early stopping after 7 epochs without improvement. We set the embedding size of XLSR to 144 and $N$ of Mamba blocks to 12 as the default configuration. The Conv1D layers in BiMamba use a kernel size of 3 and the feature dimension $D$ is set to 256.
@@ V Results and Analysis
Table I shows that DuaBiMamba performs the best on the ASVspoof 2021 LA set, achieving the lowest EER of 0.93% and min t-DCF. Its dual-column structure, with separate forward and backward Mamba layers, is beneficial for identifying fine-grained artifacts in the spoof data. However, InnBiMamba achieves the lowest EER of 1.82% on the DF set. This may be attributed to its shared input and output projection layers, which indicate that the architecture is better suited to handling the compression-induced distortions characteristic of DF data. ExtBiMamba, while providing separate pathways, may not capture the interaction between forward and backward dependencies as effectively as DuaBiMamba’s concatenation approach. It is also noted that the performance declines without BiMambas as shown in the first row of Table I, indicating its critical role. As DuaBiMamba is better at detecting diverse artifacts of LA and DF data collectively, we propose to use it as the backbone for the XLSR-Mamba model.
@@ V-D Evaluation on In-the-Wild dataset
We also evaluate our proposed model with the In-the-Wild dataset which is more challenging and capable of reflecting the generality of models due to similar background noise, emotional tone, and duration across spoofed and genuine instances. Table III shows that the proposed XLSR-Mamba model achieves the lowest EER of 6.71%, outperforming other SOTA single systems on this challenging dataset. The bidirectional state space modeling inherent in XLSR-Mamba allows for better capture of long-range dependencies. XLSR-Mamba’s structure enhances its sensitivity to long-range timing and rhythm inconsistencies, which are often subtle in deepfake speech and essential for accurately distinguishing between real and spoofed samples in realistic scenarios.